# Palpation Alters Auditory Pain Expressions with Gender-Specific Variations in Robopatients

### v1

@@ IV Experiment
Five healthy participants (2 males, 3 females) aged between 25 and 46 years (mean-34.5, SD-11.77) from the University of Cambridge, with no prior experience in pain to palpation mapping and no history of sensorimotor impairments were chosen to participate in this experiment.
@@ 5. Which attribute you paid attention to first: facial expression or pain sound or both? Pain sound
Figure 5: A) The force distribution for a given attempt for palpation is plotted against time. This corresponds to the target force of 15N. B) Box and whisker plots indicate the spread of actual palpation forces observed for each target force for the robot’s pain expressions of each gender
@@ V Results and Discussions
Fig. 5A indicates the force distribution during palpation ( period within 5 s) in a single trial. The target force level is marked. Fig. 5B illustrates the variance between the target and recorded maximum force in each trial for both male (p=4.54e-136 $<0.05$) and female (p=5.88e-164 $<0.05$) pain expressions. While there were variations in the range of forces we could observe for each target force level, average actual forces converged around 40 and 60 N for male and female pain expressions, respectively. These trends could have often been influenced by participants’ subjective comfort.
Fig. 6 shows the cumulative probability trends considering 120 trials per gender per participant, which indicates the adaptability of participants in learning pain expressions across different conditions. The early trials demonstrate variability, a hallmark of the system’s exploration phase, where different combinations of pain sounds and expressions are tested. As the trials progress, PPO’s optimisation process leads to increased stability and consistency, with participants’ preferences becoming more predictable and the system converging toward higher probabilities of selecting optimal responses. PPO’s ability to adapt in real-time is reflected in the relatively steady learning curves. These learning curves show how participants’ feedback is incorporated into the system to improve the accuracy of selected pain expressions. In addition, relatively smoother trajectories in some participants from the beginning compared to others may be explained by individual differences in feedback interpretation. One key insight is that the variability in pain expression among participants highlights the importance of having a flexible, adaptive learning model. Some individuals may show little reaction to discomfort due to high pain tolerance, while others may react strongly even to minor stimuli. PPO’s ability to adjust to these individual differences is reflected in the distinct learning curves of participants. These show that the system can handle varying feedback with sample efficiency. PPO’s capacity to adjust policies dynamically allows the system to account for these variances, resulting in distinct learning trajectories for different individuals.
@@
We plotted all the graphs separately for pain expressions of each gender to observe the trends in participants towards different genders. Fig. 7A and B show that the actual forces for all target forces average around 40 N for all amplitude levels (p=1.421e-5$<0.05$) of female pain sounds we selected throughout the experiment. This trend existed for male pain expressions as well, but the palpation force averaged around 60 N. Interestingly, this trend persisted for different pitch levels (p=1.421e-5$<0.05$) as well. This can be seen in Fig. 7C and D. However, when the actual force for each palpation was considered for the last 20% of all attempts, we could see a variation in this pattern. This can be seen in Fig. 7E-H. For all types of pain expressions (both male and female, both pitch (p=0.374 $>0.05$) and amplitude (p=0.374 $>0.05$)), the actual force averaged approximately 40 N for different pitch and amplitude levels. We consider this a result of the learning process and the tendency for individuals to settle on a comfortable way to palpate over time.
@@
Results of the experiment conclude that there is a complex and multidimensional relationship between palpation force and auditory pain expressions. While general trends were observed—such as lower amplitudes being preferred for lower forces and higher amplitudes for higher forces—participants’ subjective comfort often influenced their feedback. This introduces potential bias in the mapping of pain sounds, indicating that reverse correlation may complicate the alignment between the objective experience of pain and the robot’s expressions. The study highlights the challenges in accurately generating pain sounds. It suggests that adaptive, real-time learning models may be more effective than relying on predefined datasets for pain expression generation.
@@ V-A Post Study Questionnaire
While the study yielded positive findings, several limitations exist. The pain-to-sound mapping was limited to three levels, making it difficult for participants to differentiate feedback at maximum pain intensity. Additionally, the small sample size of naive participants may have reduced the statistical power, resulting in insignificant outcomes. This study is presented as a pilot to demonstrate how RL can make emotion-based human experiments less time consuming and more tailored for individual differences. This approach supports robot learning as well as the integration of vocal pain feedback into robopatient for abdominal palpation training. We plan to expand our investigations with a larger number of participants from diverse genders and ethnic backgrounds and participants with different palpation experience levels to address these limitations. In addition, we intend to expand the experiments to medical professionals and patients with different types of pain as well.
V-B Learning Perspective and Challenges
@@
Two main challenges in adaptive experimental design are, obtaining the accuracy of the prediction model and the large search space of possible experiments. If the prediction model is not accurate, then the estimates of information gained, and experimental utility would be affected. Hence accurate and well trained and calibrated prediction models are important. In many applications, the set of possible experimental designs is very large- hence need better ways to efficiently search the design space using ML.
@@ VI CONCLUSIONS
Our study highlights how machine learning can be used by a robot to dynamically adjust auditory pain expressions based on palpation force. This optimises responses based on palpation effort and feedback, which can be adopted to co-optimise human responses in human studies in general. The results of the experiment conclude that there is a complex, multidimensional relationship between palpation force and auditory pain expressions. General trends were observed—such as the existence of gender differences in pain vocalisations, where a lower pitch and consistent amplitude generally characterised male pain sounds. In comparison, female expressions showed higher pitch and more variable intensity. This introduces potential bias in the mapping of pain sounds, indicating that reverse correlation may complicate the alignment between the objective experience of pain and the robot’s expressions. The study highlights the challenges in accurately generating pain sounds and suggests that adaptive, real-time learning models may be more effective than relying on predefined datasets for pain expression generation. Despite the limitations, this study lays the groundwork for using reinforcement learning to personalise pain-palpation mapping and improve the accuracy of robopatient simulations in medical training. Moreover, our approach could be a cornerstone study for mapping the relationships between other feedback modalities such as visuo-haptic, visuo-auditory and many other combinations. Further studies should aim to incorporate diverse populations to generalise the findings and refine the accuracy of palpation-pain sound mapping that can create a significant impact on high-fidelity physical twins. While this study utilized PPO, we acknowledge that alternative RL methods have different trade-offs. PPO was selected due to its balance of computational efficiency, policy stability, and real-time adaptability in human-robot interaction settings. Future studies could conduct direct experimental comparisons to validate the theoretical advantages discussed in this work.

### v2

@@ IV Experiment
Five healthy participants (2 males, 3 females) aged between 25 and 46 years (mean-34.5, SD-11.77) from the University of Cambridge, with no prior experience in pain to palpation mapping and no history of sensorimotor impairments were chosen to participate in this experiment. Participants were chosen so they represent diverse backgrounds such as ethnic (European-2, Asian-2, and American-1), gender (male, female), and technical and non-technical experience to capture a wide range of perceptual responses and interaction styles. This diversity introduces natural variability into the system, helping to evaluate the generalizability of the proposed approach and simulate realistic user scenarios during the experiment.
@@ 5. Which attribute you paid attention to first: facial expression or pain sound or both? Pain sound
Figure 5: A) The force distribution for a given attempt for palpation is plotted against time. This corresponds to the target force of 15N. B) Box and whisker plots indicate the spread of actual palpation forces observed for each target force for the robot’s pain expressions of each gender.
@@ V Results and Discussions
Fig. 5A indicates the force distribution during palpation ( period within 5 s) in a single trial. The target force level is marked. Fig. 5B illustrates the variance between the target and recorded maximum force in each trial for both male (p=4.54e-136 <0.05<0.05) and female (p=5.88e-164 <0.05<0.05) pain expressions. While there were variations in the range of forces we could observe for each target force level, average actual forces converged around 40 and 60 N for male and female pain expressions, respectively. These trends could have often been influenced by participants’ subjective comfort.
Fig. 6 shows the cumulative probability trends considering 120 trials per gender per participant, which indicates the adaptability of participants in learning pain expressions across different conditions. The early trials demonstrate variability, a hallmark of the system’s exploration phase, where different combinations of pain sounds and expressions are tested. As the trials progress, PPO’s optimization process leads to increased stability and consistency, with participants’ preferences becoming more predictable and the system converging toward higher probabilities of selecting optimal responses. PPO’s ability to adapt in real-time is reflected in the relatively steady learning curves. These learning curves show how participants’ feedback is incorporated into the system to improve the accuracy of selected pain expressions. In addition, relatively smoother trajectories in some participants from the beginning compared to others may be explained by individual differences in feedback interpretation. One key insight is that the variability in pain expression among participants highlights the importance of having a flexible, adaptive learning model. Some individuals may show little reaction to discomfort due to high pain tolerance, while others may react strongly even to minor stimuli. PPO’s ability to adjust to these individual differences is reflected in the distinct learning curves of participants. These show that the system can handle varying feedback with sample efficiency. PPO’s capacity to adjust policies dynamically allows the system to account for these variances, resulting in distinct learning trajectories for different individuals.
@@
We analyzed pain expression responses separately for male and female vocalizations to examine participant trends across genders. As shown in Fig.7A and B, the actual palpation forces for all target force levels averaged around 40 N when participants responded to female pain sounds (p=1.421e-5 <<0.05). A similar pattern was observed for male pain sounds (Fig.7C and D), but with a higher average palpation force of approximately 60 N, suggesting that participants tolerated or applied greater pressure when hearing male pain expressions. Notably, this trend persisted across both pitch and amplitude variations.
However, when focusing on the final 20% of palpation attempts (Fig. 7E–H), the distinction between male and female pain sound responses diminished. In this phase, participants converged on an average force of around 40 N across all conditions (p=0.374>>0.05), regardless of gender or acoustic parameters. We interpret this convergence as part of a learning or calibration process, where participants adjusted their behavior to a self-determined comfort zone over time.
The observed gender-based variations in initial responses may be explained by both physiological and psychosocial factors. Prior research suggests that emotional salience and perceived vulnerability differ based on voice gender; for example, female vocalizations may be interpreted as more urgent or emotionally expressive [37], leading to more cautious palpation. In contrast, male voices may be perceived as more tolerant or stoic, resulting in participants exerting higher forces. These findings align with literature in social cognition and gendered perception of pain, which suggests that listener biases and social norms can influence perceived severity and response behavior [42, 47].
These results underscore the need for personalized and adaptive pain models in robopatient systems, especially when integrating human-in-the-loop feedback. Future work will explore individual and cultural biases more deeply and examine how robot voice characteristics modulate user behavior in procedural tasks.
@@
Results of the experiment conclude that there is a complex and multidimensional relationship between palpation force and auditory pain expressions. While general trends were observed—such as lower amplitudes being preferred for lower forces and higher amplitudes for higher forces—participants’ subjective comfort often influenced their feedback. This introduces potential bias in the mapping of pain sounds, indicating that reverse correlation may complicate the alignment between the objective experience of pain and the robot’s expressions. The study highlights the challenges in accurately generating pain sounds. It suggests that adaptive, real-time learning models may be more effective than relying on predefined datasets for pain expression generation. Hence, co-optimizing with machine learning may offer a more adaptable approach to studying multidimensional human data.
Our experimental results revealed gender-related differences in the perception and expression of pain sounds, particularly in the thresholds at which pain was perceived as intense. While this study did not aim to investigate these differences in depth, prior literature suggests that both physiological and psychosocial factors may contribute.
On the physiological side, differences in pain sensitivity, hormonal modulation (e.g., estrogen’s influence on nociceptive processing), and auditory perception thresholds have been reported across genders [10]. Psychologically, gender norms and expressivity may influence how participants perceive and evaluate pain sounds—e.g., males may downregulate perceived intensity, while females may be more attuned to emotional vocalizations [47, 42]. Such variations may persist across gender, age, and ethnicity. However, broader observations involving diverse participants and robotic systems will be necessary to generalize these findings within the domain of human–robot interaction.
@@ V-A Post Study Questionnaire
V-B Limitations, Learning Perspective and Challenges
While the study yielded positive findings, several limitations exist. The pain-to-sound mapping was limited to three levels, making it difficult for participants to differentiate feedback at maximum pain intensity. Additionally, the small sample size of naive participants may have reduced the statistical power, resulting in insignificant outcomes. This study is presented as a pilot to demonstrate how ML can make emotion-based human experiments less time consuming and more tailored for individual differences. This approach supports robot learning as well as the integration of vocal pain feedback into robopatients for abdominal palpation training. We plan to expand our investigations with a larger number of participants from diverse genders and ethnic backgrounds and participants with different palpation experience levels to address these limitations. In addition, we intend to expand the experiments to medical professionals and patients with different types of pain as well.
@@ V-B Limitations, Learning Perspective and Challenges
Two main challenges in adaptive experimental design are, obtaining the accuracy of the prediction model and the large search space of possible experiments. If the prediction model is not accurate, then the estimates of information gained, and experimental utility would be affected. Hence accurate and well trained and calibrated prediction models are important. In many applications, the set of possible experimental designs is very large- hence need better ways to efficiently search the design space using ML. High dimensionality in concepts such as pain could make this even more tedious.
@@
While this work considers palpation using discrete force levels, real-world palpation occurs in a continuous force space. To better reflect this, we selected PPO, which is well-suited for continuous action spaces. This choice is made to facilitate future extensions toward modeling palpation as a continuous control task. Furthermore the current study does not aim to benchmark reinforcement learning algorithms, it introduces a novel paradigm for human–ML co-optimization in affective robotics. Our focus is on demonstrating feasibility and exploring trends rather than establishing algorithmic superiority. Future work will include comparative evaluations with alternative RL and data-driven methods to assess model generalizability and performance.
@@ VI CONCLUSIONS
Our study highlights how machine learning can be used by a robot to dynamically adjust auditory pain expressions based on palpation force. This optimizes responses based on palpation effort and feedback, which can be adopted to co-optimize human responses in human studies in general. The results of the experiment conclude that there is a complex, multidimensional relationship between palpation force and auditory pain expressions. General trends were observed—such as the existence of gender differences in pain vocalisations, where a lower pitch and consistent amplitude generally characterised male pain sounds. In comparison, female expressions showed higher pitch and more variable intensity. This introduces potential bias in the mapping of pain sounds, indicating that reverse correlation may complicate the alignment between the objective experience of pain and the robot’s expressions. The study highlights the challenges in accurately generating pain sounds and suggests that adaptive, real-time learning models may be more effective than relying on predefined datasets for pain expression generation. Despite the limitations, this study lays the groundwork for using reinforcement learning to personalise pain-palpation mapping and improve the accuracy of robopatient simulations in medical training. Moreover, our approach could be a cornerstone study for mapping the relationships between other feedback modalities such as visuo-haptic, visuo-auditory and many other combinations. Further studies should aim to incorporate diverse populations to generalise the findings and refine the accuracy of palpation-pain sound mapping that can create a significant impact on high-fidelity physical twins. While PPO provide ideal for out interaction constraints, future work could benchmark newer RL methods in scaled deployments with richer feedback modalities.